# Recent Advances of Foundation Language Models-based Continual Learning: A Survey

### v1

@@ 4.3.1. PLMs-based CIL
The study presented in (Monaikul et al., 2021) introduces ExtendNER, a novel framework for continual learning in Named Entity Recognition (NER) that obviates the need for extensive re-annotation. This framework employs a knowledge distillation (KD) technique, wherein a pre-existing named entity recognition (NER) model, termed the ”teacher”, imparts knowledge to a newly developed model, termed the ”student”. The student model, designed to identify new entity types, progressively learns by emulating the teacher model’s responses to a new dataset. This method allows the student to acquire the ability to recognize new entities while retaining knowledge of previously identified ones. Additionally, Liu et al. (Liu et al., 2019b) propose an innovative method for learning distributed representations of sentences. This method initiates with the configuration of sentence encoders using features independent of any specific corpus, followed by iterative updates through Boolean operations on conceptor matrices. This technique ensures that the encoders maintain their performance on existing datasets while adapting effectively to new data.
(a) IDBR
@@
Figure 7. Frameworks in CIL: IDBR (PLM-based) (Huang et al., 2021), PLE (PLM-based) (Li et al., 2022b), Adaptation-CLIP (VLM-based) (Liu et al., 2023a).
Huang et al. (Huang et al., 2021) introduced an innovative methodology known as Information Disentanglement-based Regularization (IDBR) to address the enduring challenges associated with continual text classification. This method effectively disentangles the hidden spaces of text into task-generic and task-specific representations, employing distinct regularization strategies to enhance knowledge retention and facilitate generalization. Furthermore, the integration of two auxiliary tasks, namely next sentence prediction and task-id prediction, serves to augment the learning process by reinforcing the separation between generic and specific representational spaces.
@@
Introduced by Varshney et al. (Varshney et al., 2022), Prompt Augmented Generative Replay (PAGeR) is a method that enables continual learning in intent detection without retaining past data. PAGeR leverages pre-trained language models to generate intent-specific utterances specific to new intents while preserving accuracy on existing ones. Unlike exemplar replay, which stores specific examples, PAGeR is structured to selectively maintain relevant contexts for each specified intent, which are then employed as generation prompts. This approach combines utterance generation and classification into one model, enhancing knowledge transfer and optimization.
@@
In the domain of task-oriented dialogue systems, Continual Few-shot Intent Detection (CFID) focuses on recognizing new intents with few examples. Li et al. (Li et al., 2022b) propose a Prefix-guided Lightweight Encoder (PLE) to address this by using a parameter-efficient tuning method that combines a Continual Adapter module with a frozen Pre-trained Language Model (PLM) and a Prefix-guided Attention mechanism to reduce forgetting. To further mitigate forgetting, the Pseudo Samples Replay (PSR) strategy reinforces prior knowledge by replaying crucial samples from previous tasks. The Teacher Knowledge Transfer (TKT) strategy uses distillation to transfer task-specific knowledge to maintain performance on new tasks. Additionally, the Dynamic Weighting Replay (DWR) strategy dynamically adjusts weights of previous tasks to balance new knowledge acquisition with the revision of old tasks, navigating the variability and potential negative impacts of prior tasks
The Efficient Parameter Isolation (EPI) method, introduced in Wang et al. (Wang et al., 2023c), assigns unique subsets of private parameters to each task alongside a shared pre-trained model. This approach ensures precise parameter retrieval and has been shown to outperform non-caching methods in continual language learning benchmarks, while remaining competitive with caching methods. Furthermore, EPI employs random static masking to reduce storage requirements, increasing its viability in resource-constrained environments.
In complex system environments, efficient and adaptable machine learning architectures are crucial, especially for classification tasks with sequentially presented data. Wójcik et al. (Wójcik et al., 2023) devise a novel architecture, Domain and Expertise Ensemble (DE&E), comprising a feature extractor, classifier, and gating mechanism. The feature extractor employs a multi-layer neural network to convert input data into embeddings, while the classifier, a mixture of binary class-specific experts, leverages a gating mechanism to select the appropriate expert for the current input dynamically. This Mixture of Experts-based method promotes incremental learning by training experts with class-specific samples and combines their outputs during testing to derive the final classification.
@@ 4.3.2. VLMs-based CIL
Kim et al. (Kim et al., 2024) introduce VLM-PL, a novel approach for class incremental object detection that incorporates new object classes into a detection model without forgetting old ones. Utilizing a visual-language model (VLM), this method enhances the pseudo-labeling process to improve the accuracy and performance of object detection in continual learning settings. VLM-PL starts with a pre-trained detector to generate initial pseudo-labels, which are then validated through a visual-language evaluation using a specially designed hint template. Accurate pseudo-labels are retained and combined with ground truth labels to train the model, ensuring it remains proficient in recognizing both new and previously learned categories. Cao et al. (Cao et al., 2024) introduce a framework for a Generative Multi-modal Model (GMM) that leverages large language models for class-incremental learning. This innovative approach entails the generation of labels for images by employing an adapted generative model. Following the production of detailed textual descriptions, a text encoder is utilized to extract salient features from these descriptions. These extracted features are subsequently aligned with existing labels to ascertain the most fitting label for classification predictions.
PROOF (Zhou et al., 2023c) develops a method to enhance model memory retention when adapting to downstream tasks. This method involves a projection technique that maps pre-trained features into a new feature space designed to preserve prior knowledge. Additionally, to effectively utilize cross-modal information, PROOF introduces a fusion module that employs an attention mechanism. This module adjusts both visual and textual features simultaneously, enabling the capture of semantic information with enhanced expressive power. Recently, Jha et al. (Jha et al., 2024) present a new approach for adapting pre-trained vision-language models like CLIP to new tasks without forgetting previous knowledge. It employs a Variational Inference framework to probabilistically model the distribution of visual-guided text features, enhancing fine-tuning reliability by accounting for uncertainties in visual-textual interactions. Key to CLAP is the visual-guided attention (VGA) module, which aligns text and image features to prevent catastrophic forgetting. Additionally, CLAP includes lightweight, task-specific inference modules that learn unique stochastic factors for each task, allowing continuous adaptation and knowledge retention.
@@
Liu et al. (Liu et al., 2023a) introduce Adaptation-CLIP, which employs three strategies for CLIP’s continual learning: linear adapter, self-attention adapter, and prompt tuning. The first two strategies add a linear layer and a self-attention mechanism, respectively, after the image encoder while freezing the remaining architecture. The third, prompt tuning, integrates trained prompts into the text encoder to enhance task comprehension and splices these with prior prompts to maintain continuity. To prevent catastrophic forgetting, a parameter retention strategy preserves significantly altered parameters from $M_{t-1}$ to $M_{t}$, ensuring stability and effective continual learning.
@@
Khan et al. (Khan et al., 2023) introduce two notable advancements: an enhanced prompt pool key query mechanism and category-level language guidance. The key query mechanism uses CLS features to improve prompt selection, featuring key replacement with a fixed CLS tag and dynamic mapping to task-level language representations, thereby enhancing accuracy and robustness across different tasks. Meanwhile, category-level language guidance is implemented in the vision transformer to better align output features with category-specific language representations, significantly improving task handling and category differentiation, leading to improved model performance.
@@ 5.1.1. PLMs-based HTB
The Hard Task Boundary (HTB) setting has been developed to enable continuous knowledge acquisition by learning models from a dynamically changing stream of textual data, without the need for dataset identifiers. For example, Shen et al. (Shen et al., 2019) have implemented HTB in slot filling, Michieli et al. (Michieli et al., 2023) have utilized it in audio classification, and Vander et al. (Vander Eeckt et al., 2023) have explored its use in automatic speech recognition.
@@
Continual learning (CL) methodologies, particularly pertinent to online scenarios, encompass a variety of approaches. These include parameter-isolation-based methods (de Masson D’Autume et al., 2019; Wang et al., 2020), replay-based methods (Holla et al., 2020) and regularization-based methods (Vander Eeckt et al., 2023; Liu et al., 2021b). MBPA++ (de Masson D’Autume et al., 2019) introduces a framework for lifelong language learning, enabling a pre-trained model to learn continually from textual examples without requiring labeled datasets. It employs an episodic memory system with sparse experience replay and local adaptation techniques to prevent catastrophic forgetting. Extending this framework, Meta-MBPA++ (Wang et al., 2020) integrates three core lifelong learning principles, enhancing performance in text classification and question-answering tasks while using only 1% of the typical memory usage. Liu et al. (Liu et al., 2021b) introduce a regularization-based strategy, CID, for lifelong intent detection. This method uses cosine normalization, hierarchical knowledge distillation, and inter-class margin loss to tackle the challenges of data imbalances in the lifelong intent detection (LID) task, aiming to mitigate the negative impacts associated with these imbalances.
@@
(b) CBA
Figure 8. Frameworks in Online Continual Learning: MBPA++ (PLM-based HTB/BTB) (de Masson D’Autume et al., 2019), CBA (VLM-base BTB) (Wang et al., 2023d).
@@ 5.2.1. PLMs-based BTB
MBPA++ (de Masson D’Autume et al., 2019) and Meta-MBPA++ (Wang et al., 2020) can also exemplify models capable of adapting to environments with indistinct task boundaries. TPEM (Geng et al., 2021) employs a tripartite approach within an encoder-decoder framework, consisting of pruning, expanding, and masking techniques. Pruning helps preserve essential information from previous tasks, expansion increases the model’s capacity to accommodate new tasks, and masking reduces interference from previous tasks’ weights, thus enhancing learning efficiency. Originally, online meta-learning (OML) (Javed and White, 2019) and a neuromodulatory meta-learning algorithm (ANML) (Beaulieu et al., 2020) were intended to continuously learn new sequences of tasks during the testing phase. Holla et al. (Holla et al., 2020) adapt these algorithms for a conventional continual learning context, where the focus is on evaluating performance on previously encountered tasks. The enhanced versions, named OML-ER and ANML-ER, incorporate an episodic memory module designed for experience replay.
@@ 5.2.2. VLMs-based BTB
Traditional Methods.
Cui et al. (Cui et al., 2024) propose the Dynamic Knowledge Rectification (DKR) framework, designed to mitigate the propagation of incorrect information in foundation LMs. The DKR framework operates by initially leveraging an existing model to identify and exclude obsolete or erroneous knowledge when confronted with new data. Subsequently, a rectification process is employed to amend these inaccuracies while ensuring the preservation of valid data associations. This process is especially vital when integrating new data, as it prevents the perpetuation of outdated or incorrect information. In cases where data is inaccessible through existing models, DKR utilizes paired ground-truth labels to support the continuous evolution of knowledge bases, thereby enhancing the model’s accuracy.
Parameter-Efficient Tuning Methods.
Wang et al. (Wang et al., 2023d) introduce the Continual Bias Adaptor (CBA) (Figure 8(b)), a novel method designed to enhance the efficacy of online CL by mitigating catastrophic forgetting. The CBA method dynamically modifies the classifier’s network to adapt to significant shifts in data distribution observed during training, thereby preserving knowledge from earlier tasks. Notably, the CBA module can be deactivated during testing phases, eliminating additional computational burdens or memory demands. This feature highlights the practicality and applicability of the CBA in real-world scenarios.
Instruction Tuning-based Methods.
In the BTB scenario, the Mask and Visual Prompt tuning (MVP) method, as detailed by Moon et al. (Moon et al., 2023), addresses challenges like intra- and inter-task forgetting and class imbalance effectively. MVP features instance-wise logit masking to prevent irrelevant information retention, contrastive visual prompt tuning loss to ensure consistent prompt selection, gradient similarity-based focal loss to focus on overlooked samples, and adaptive feature scaling to balance the integration of new knowledge with existing data retention.
Online Learning for CV
Recent advancements in class-incremental online learning predominantly address computer vision tasks, encapsulating various methodologies such as regularization-based (Yin et al., 2021; Gu et al., 2022; Guo et al., 2022; Bang et al., 2022), replay-based (Aljundi et al., 2019; Shim et al., 2021; Caccia et al., 2021; Lin et al., 2023; Zhang et al., 2022a), distillation-based (He et al., 2020; Fini et al., 2020; Koh et al., 2022), and gradient-based approaches (Jin et al., 2021b; Guo et al., 2023; Chen et al., 2020; Yoon et al., 2021; Ororbia et al., 2022). Among these, Koh et al. (Koh et al., 2021) introduced the Class-Incremental Blurry (CLIB) model, which distinguishes itself through its task-free nature, adaptability to class increments, and prompt response to inference queries, showing superior performance over traditional continual learning methods. In a related vein, Gunasekara et al. (Gunasekara et al., 2023) explore the Online Streaming Continual Learning (OSCL), a hybrid of Stream Learning (SL) and Online Continual Learning (OCL), which integrates aspects of both domains. Furthermore, issues of shortcut learning and bias in online continual learning have been tackled by Wei et al. (Wei et al., 2023) and Chrysalis et al. (Chrysakis and Moens, 2023). Additionally, Semola et al. (Semola et al., 2022) propose Continual-Learning-as-a-Service (CLaaS), a service model that leverages continual learning to monitor shifts in data distribution and update models efficiently. This array of developments highlights the dynamic capabilities of online continual learning frameworks to effectively address complex challenges across a variety of computer vision tasks.
Table 1. The statistics information of the existing CL datasets. #D/T/C means the number of domains/tasks/classes, respectively.
Datasets #Train #Val #Test #Total CL Settings NLP Problems Language #D/T/C
@@
Progressive Prompts (Razdaibiedina et al., 2023) - - - - TIL Sentiment analysis, topic classification, boolean QA, QA, paraphrase detection, word sense disambiguation, natural language inference English 15 tasks
MeLL (Wang et al., 2021a) 1,430,880 173,781 118,240 1,722,901 TIL Intent classification English 1184 tasks
Continual-T0 (Scialom et al., 2022b) 800,000 - 33,382 833,382 TIL Text Simplification, Headline Generation with Constraint, Haiku Generation, Covid QA, Inquisitive Question Generation, Empathetic Dialogue Generation, Explanation Generation, Twitter Stylometry English 8 tasks
COPR (Zhang et al., 2023b) - - - - TIL QA tasks, summary task, positive file review generation task English 3 tasks
Adaptive Compositional Modules (Zhang et al., 2022b) 50,725 - 27,944 78,669 TIL Natural language generation, SQL query generation, Summarization and Task-oriented dialogue English 8 tasks
CODETASKCL (Yadav et al., 2023) 181,000 9,700 10,000 200,700 TIL Code generation, code translation, code summarization, and code refinement Hybrid 4 tasks
Lifelong Simple Questions (Wang et al., 2019b) - - - - TIL single-relation questions English 20 tasks
Lifelong FewRel (Wang et al., 2019b) - - - - TIL few-shot relation detection English 10 tasks
InstrDialog (Zhang et al., 2023a) 9,500 950 1,900 12,350 TIL Dialogue state tracking, dialogue generation, intent identification English 19 tasks
InstrDialog++ (Zhang et al., 2023a) 3,800 1,900 3,800 9,500 TIL Dialogue Generation, Intent Identification, Dialogue State Tracking, Style Transfer, Sentence Ordering, Word Semantics, Text Categorization, Pos Tagging, Fill in The Blank, Program Execution, Question Generation, Misc, Coherence Classification, Question Answering, Summarization, Commonsense Classification, Wrong Candidate Generation and Toxic Language Detection English 38 tasks
ConTinTin (Yin et al., 2022) - - - - TIL Question generation tasks (QG), answer generation tasks (AG), classification tasks (CF), incorrect answer generation tasks (IAG), minimal modification tasks (MM) and verification tasks (VF) English 61 tasks
Tencent TL (Yuan et al., 2021) - - - - TIL personalized recommendations and profile predictions English 6 tasks
Movielens (Yuan et al., 2021) - - - - TIL personalized recommendations and profile predictions English 3 tasks
@@
NAVER Shopping (Kim et al., 2023) - - - - TIL search query prediction tasks English 6 tasks
TRACE (Wang et al., 2024a) 40,000 - 16,000 56,000 TIL Domain-specific task, Multi-lingual task, Code completion task, Mathematical reasoning task Hybrid 8 tasks
@@
ABSC (Ke and Xu, 2021) 3,452 150 1,120 4,722 DIL Aspect-based sentiment classification English 19 Domains
DecaNLP (Sun et al., 2020a) 169,824 - 32,116 201,940 DIL Question answering, semantic parsing, sentiment analysis, semantic role labeling, and goal-oriented dialogue English 5 domains
Foundational text classification (Sun et al., 2020a) 115,000 - 7,600 122,600 DIL News classification, sentiment analysis, Wikipedia article classification, and question-and-answer categorization English 5 domains
RVAE_LAMOL (Wang et al., 2022a) 15,870 - 5,668 21,538 DIL Oriented dialogue of the restaurant reservation task, semantic role labeling, sentiment classification English 3 domains
COPR (Zhang et al., 2023b) - - - - DIL - English 18 domains
@@
SGD (Zhu et al., 2022) 38,745 5,210 11,349 40,287 DIL Dialogue state tracking English 19 Domains
@@
CPT (Ke et al., 2022) 3,121,926 - - 3,121,926 DIL Domain-adaptive pre-training task English 4 Domains
@@
CKL(Jang et al., 2022b) - - - 30,372 DIL Domain-adaptive pre-training task English 3 Domains
@@
ELLE (Qin et al., 2022) - - - - DIL Domain-adaptive pre-training task English 5 Domains
@@
Domain-incremental Paper Stream (Jin et al., 2022) - - - - DIL Relation extraction and named entity recognition English 4 domains
Chronologically-ordered Tweet Stream (Jin et al., 2022) - - - - DIL multi-label hashtag prediction and single-label emoji prediction English 4 domains
AdapterCL (Madotto et al., 2020) 31,426 4,043 4,818 40,287 DIL Intent classification, Dialogue State Tracking (DST), Natural Language Generation (NLG), end-to-end (E2E) modeling English 37 Domains
@@
DE&E (Wójcik et al., 2023) 28,982 - 12,089 41,071 CIL Text classification English 3 tasks
EPI (Wang et al., 2023c) 12,840 3,524 6,917 23,281 CIL Text classification, topic classification English 13 Classes
@@
PAGeR (Varshney et al., 2022) 59,754 7,115 15,304 82,173 CIL Intent classification English 355 Classes
PLE (Li et al., 2022b) 4,669 4,650 31,642 40,961 CIL Intent classification English 477 Classes
@@
CoNLL-03 (Monaikul et al., 2021) 23,326 5,902 5,613 34,841 CIL Named Entity Recognition English 4 Classes
OntoNotes (Monaikul et al., 2021) 107,169 16,815 10,464 134,448 CIL Named Entity Recognition English 6 Classes

### v2

@@ 5.3.1. PLMs-based CIL
ExtendNER (Monaikul et al., 2021) (Figure 7(a)) introduces a continual learning framework for Named Entity Recognition (NER) that reduces re-annotation efforts. It uses knowledge distillation, where a ”teacher” NER model helps a ”student” model learn to recognize new entity types while retaining knowledge of previous ones. SRC (Liu et al., 2019b) focuses on sentence representation learning by initializing corpus-independent encoders and refining them through Boolean operations on conceptor matrices. IDBR (Huang et al., 2021) improves continual text classification by disentangling task-generic and task-specific representations. It uses regularization techniques and auxiliary tasks like next sentence prediction to enhance knowledge retention and generalization, outperforming many existing methods across various task sequences.
(a) ExtendNER
@@
Figure 7. Frameworks in CIL: ExtendNER (PLM-based) (Monaikul et al., 2021), PLE (PLM-based) (Li et al., 2022b), Adaptation-CLIP (VLM-based) (Liu et al., 2023a).
@@
Instead of retaining examples, PAGeR (Varshney et al., 2022) uses PLMs to generate intent-specific utterances for new tasks while maintaining performance on previous ones. It selectively preserves relevant contexts as prompts, ensuring efficient memory usage. By utilizing PLMs, PAGeR is effective in real-world scenarios where new intents and relations are continuously introduced, offering a memory-efficient alternative to traditional methods.
@@
In task-oriented dialogue systems, continuous few-shot intent detection (CFID) focuses on recognizing new intents with a few examples. To address this, PLE (Li et al., 2022b) (Figure 7(b)) employs a parameter-efficient tuning approach that integrates a Continual Adapter module with a frozen PLM and a Prefix-guided Attention mechanism. This design effectively reduces forgetting while minimizing the number of trainable and stored parameters, making it an ideal choice for real-world applications where computational efficiency is critical. Similarly, EPI (Wang et al., 2023c) allocates task-specific private parameters alongside a shared model, ensuring precise retrieval and superior performance in continual learning. It reduces storage needs through random static masking and consistently outperforms rehearsal-free methods, even competing with rehearsal-based ones on far-domain datasets. For environments with sequentially presented data, DE&E (Wójcik et al., 2023) uses a mixture of binary class-specific experts, dynamically selecting the best expert for each input, facilitating incremental learning by combining expert outputs for final classification.
@@ 5.3.2. VLMs-based CIL
VLM-PL (Kim et al., 2024) leverages a VLM to enhance the pseudo-labeling process, integrating new object classes into a detection model without compromising previously learned categories. GMM (Cao et al., 2024) leverages LLMs for class-incremental learning. This innovative approach entails the generation of labels for images by employing an adapted generative model. PROOF (Zhou et al., 2023c) maps pre-trained features into a new feature space designed to preserve prior knowledge. To enhance the use of cross-modal information, it incorporates a fusion module with an attention mechanism, allowing the model to retain essential features when adapting to new tasks. CLAP (Jha et al., 2024) offers a new method for adapting VLMs to novel tasks without losing previously learned knowledge. This approach employs a Variational Inference framework to probabilistically model the distribution of visual-guided text features, enhancing fine-tuning reliability by accounting for uncertainties in visual-textual interactions.
@@
Adaptation-CLIP (Liu et al., 2023a) (Figure 7(c)) employs three strategies for continual learning: a linear adapter, a self-attention adapter, and prompt tuning. The adapters add layers to the image encoder while freezing the rest of the model, and prompt tuning enhances the text encoder by integrating new and prior prompts to maintain task comprehension. DIKI (Tang et al., 2024) prevents the disruption of pre-trained knowledge by using a residual attention mechanism that injects new knowledge into a frozen backbone, minimizing interference. A distribution-aware calibration scheme is employed to maintain the zero-shot capabilities of VLMs. RAPF (Huang et al., 2024) addresses class-incremental learning challenges using a pre-trained CLIP model with a linear adapter. It minimizes interference by identifying neighboring categories and applying hinge loss, while generating old class features from a Gaussian distribution to aid classification. STAR-Prompt (Menabue et al., 2024) introduces a two-level prompting mechanism to balance stability and plasticity in large models like CLIP and ViT, reducing catastrophic forgetting while enabling continuous learning.
@@
LGCL (Khan et al., 2023) introduces two key innovations: a refined prompt pool key query mechanism and category-level language guidance. The former leverages CLS features to optimize prompt selection with dynamic task-level mappings, improving accuracy and robustness, while the latter aligns output features with specific language representations in the vision transformer, enhancing task handling and model performance.
@@ 6.1.1. PLMs-based HTB
The Hard Task Boundary (HTB) setting has been developed to enable continuous knowledge acquisition by learning models from a dynamically changing stream of textual data, without the need for dataset identifiers. For example, ProgModel (Shen et al., 2019) have implemented HTB in slot filling, TAP-SLDA (Michieli et al., 2023) have utilized it in audio classification, and AOS (Vander Eeckt et al., 2023) have explored its use in automatic speech recognition.
@@
Continual learning (CL) methodologies, particularly pertinent to online scenarios, encompass a variety of approaches. These include parameter-isolation-based methods (de Masson D’Autume et al., 2019; Wang et al., 2020), replay-based methods (Holla et al., 2020) and regularization-based methods (Vander Eeckt et al., 2023; Liu et al., 2021b). MBPA++ (de Masson D’Autume et al., 2019) (Figure 8(a)) introduces a framework for lifelong language learning, enabling a pre-trained model to learn continually from textual examples without requiring labeled datasets. It employs an episodic memory system with sparse experience replay and local adaptation techniques to prevent catastrophic forgetting. Extending this framework, Meta-MBPA++ (Wang et al., 2020) integrates three core lifelong learning principles, enhancing performance in text classification and question-answering tasks while using only 1% of the typical memory usage. CID (Liu et al., 2021b) is devised for lifelong intent detection. This method uses cosine normalization, hierarchical knowledge distillation, and inter-class margin loss to tackle the challenges of data imbalances in the lifelong intent detection task, aiming to mitigate the negative impacts associated with these imbalances.
@@
(b) G-NoCL
Figure 8. Frameworks in Online Continual Learning: MBPA++ (PLM-based HTB/BTB) (de Masson D’Autume et al., 2019), G-NoCL (VLM-base BTB) (Seo et al., 2024).
6.1.2. VLMs-based HTB
PEGP (Qiao et al., 2024) utilizes various parameter-efficient tuning methods and an orthogonal gradient projection mechanism to prevent catastrophic forgetting in continual learning. Using backbone networks like ViT and CLIP, PEGP ensures that gradient updates during training are orthogonal to the feature subspace of previous tasks, preserving old knowledge while learning new ones.
@@ 6.2.1. PLMs-based BTB
MBPA++ (de Masson D’Autume et al., 2019) and Meta-MBPA++ (Wang et al., 2020) exemplify models capable of adapting to environments with indistinct task boundaries. TPEM (Geng et al., 2021) adopts a tripartite approach within an encoder-decoder framework, utilizing pruning, expanding, and masking techniques. Pruning preserves essential information from previous tasks, expansion increases model capacity for new tasks, and masking mitigates interference from prior task weights, thereby enhancing learning efficiency. Online meta-learning (OML) (Javed and White, 2019) and the neuromodulatory meta-learning algorithm (ANML) (Beaulieu et al., 2020) are initially designed to learn sequences of tasks during the testing phase. Holla et al. (Holla et al., 2020) adapts these algorithms for conventional continual learning, focusing on performance evaluation on previous tasks. Their enhanced versions, OML-ER and ANML-ER, incorporate an episodic memory module for experience replay. The S6 framework addresses the challenges of continual relation learning in real-world scenarios, where data arrives in a streaming manner with noisy labels and shifting distributions.
@@ 6.2.2. VLMs-based BTB
DKR (Cui et al., 2024) is devised to mitigate the propagation of incorrect information in foundation LMs. It operates by initially leveraging an existing model to identify and exclude obsolete or erroneous knowledge when confronted with new data. Subsequently, a rectification process is employed to amend these inaccuracies while preserving valid data associations. SIT (Wang et al., 2024a) addresses the challenge of asymmetry in Online Lifelong Learning (OLL) with VLMs, where the image encoder is tuned on the current batch of data, while the text encoder can access features from previously seen classes. SIT reformulates the loss function by restricting the text features in the comparison to only those present in the current batch, ensuring that both image and text feature updates remain balanced.
OLiVia-Nav (Narasimhan et al., 2024) is an innovative architecture designed for social navigation in human-centered environments like hospitals and offices, integrating VLMs with an online lifelong learning framework. The system employs a novel distillation approach, Social Context Contrastive Language Image Pre-training (SC-CLIP), to transfer the social reasoning capabilities of large VLMs into lightweight models, enabling real-time encoding of social and environmental context. Generative Name only Continual Learning (G-NoCL) (Seo et al., 2024) (Figure 8(b)) is a framework designed to overcome the challenges of manual annotation and web-scraped data in continual learning. The core of the model is the Prompt Refiner Module, which transforms basic concept names into diverse, context-rich prompts.
Table 1. The statistics information of the existing CL datasets. #D/T/C means the number of domains/tasks/classes for DIL, TIL and CIL, respectively.
Datasets #Train #Val #Test #Total CL Settings NLP Problems Source Language #D/T/C
@@
Progressive Prompts (Razdaibiedina et al., 2023) - - - - TIL Mixed 15 classification tasks AGNews, Amazon Reviews, Yelp Reviews, DBpedia, Yahoo Answers, MNLI, QQP, RTE, SST2, WiC, CB, COPA, MultiRC, BoolQ English 15 tasks
MeLL (Wang et al., 2021) 1,430,880 173,781 118,240 1,722,901 TIL Intent classification Snips, TOP semantic parsing, Facebook Multilingual Task Oriented Dataset, e-commerce dialogue system AliMe English 1184 tasks
Continual-T0 (Scialom et al., 2022a) 800,000 - 33,382 833,382 TIL Language generation task WikiAuto, Gigaword, Subreddit haiku, ELI5, Empathetic dialogues, e-SNLI, Twitter English 8 tasks
COPF (Zhang et al., 2023b) - - - - TIL Value alignment IMDB, HH-RLHF, Reddit TL, DR English 3 tasks
ACM (Zhang et al., 2022) 50,725 - 27,944 78,669 TIL Mixed 4 generation tasks E2E NLG, RNNLG, WikiSQL, CNN/DailyMail, MultiWOZ English 4 tasks
CODETASKCL (Yadav et al., 2023) 181,000 9,700 10,000 200,700 TIL Mixed 4 code-centric tasks CONCODE, CodeTrans, CodeSearchNet, BFP Hybrid 4 tasks
Lifelong SimpleQuestions (Wang et al., 2019b) - - - - TIL Relation extraction SimpleQuestions English 20 tasks
Lifelong FewRel (Wang et al., 2019b) - - - - TIL Relation extraction FewRel English 10 tasks
InstrDialog (Zhang et al., 2023a) 9,500 950 1,900 12,350 TIL Mixed 3 dialogue tasks SuperNI English 19 tasks
InstrDialog++ (Zhang et al., 2023a) 3,800 1,900 3,800 9,500 TIL Mixed 22 dialogue tasks SuperNI English 38 tasks
ConTinTin (Yin et al., 2022) - - - - TIL Mixed classification and generation tasks NATURAL-INSTRUCTIONS English 61 tasks
Conure (Yuan et al., 2021) - - - - TIL User representations Tencent TL, Movielens English 9 tasks
@@
NAVER Shopping (Kim et al., 2023) - - - - TIL Search query prediction Tencent TL, Movielens English 6 tasks
TRACE (Wang et al., 2024b) 40,000 - 16,000 56,000 TIL Mixed diverse tasks ScienceQA, FOMC, MeetingBank, C-STANCE, 20Minuten, CodeXGLUE, NumGLUE Hybrid 8 tasks
@@
ABSC (Ke and Xu, 2021) 3,452 150 1,120 4,722 DIL Aspect-based sentiment classification L5Domains, Liu3Domains, Ding9Domains, SemEval14 English 19 domains
LAMOL (Sun et al., 2020a) 284,824 - 39,716 324,540 DIL Mixed classification and labeling tasks AGNews, Amazon Reviews, Yelp Reviews, DBpedia, Yahoo Answers, SQuAD, WikiSQL, SST, QA-SRL, WOZ English 5 domains
RVAE_LAMOL (Wang et al., 2022a) 15,870 - 5,668 21,538 DIL Generation tasks WOZ, QA-SRL, SST English 3 domains
COPF (Zhang et al., 2023b) - - - - DIL Value alignment Standard Human Preference (SHP) English 18 domains
@@
C-PT (Zhu et al., 2022) 38,745 5,210 11,349 40,287 DIL Dialogue state tracking Schema-Guided Dialog dataset (SGD) English 19 domains
@@
CPT (Ke et al., 2022) 3,121,926 - - 3,121,926 DIL Domain-adaptive pre-training task Yelp Restaurant, AI Papers, ACL Papers, AGNews, SemEval-res, ACL-ARC, SCIERC English 4 domains
@@
CKL(Jang et al., 2022b) - - - 30,372 DIL Domain-adaptive pre-training task INVARIANTLAMA, UPDATEDLAMA, NEWLAMA, NEWLAMA-EASY English 3 domains
@@
ELLE (Qin et al., 2022) - - - - DIL Domain-adaptive pre-training task BOOKCORPUS(WB), NEWS ARTICLES(NS), AMAZON REVIEWS(REV), BIOMEDICAL PAPERS(BIO), COMPUTER SCIENCE PAPERS(CS) English 5 domains
@@
Domain-incremental Paper Stream (Jin et al., 2022) - - - - DIL Mixed 2 information extraction tasks S2ORC English 4 domains
Chronologically-ordered Tweet Stream (Jin et al., 2022) - - - - DIL Mixed 2 classification tasks Tweets English 4 domains
AdapterCL (Madotto et al., 2020) 31,426 4,043 4,818 40,287 DIL Mixed 4 dialogue tasks TaskMaster2019(TM19), TaskMaster2020(TM20), Schema Guided Dialogue(SGD), MultiWoZ English 37 domains
@@
DE&E (Wójcik et al., 2023) 28,982 - 12,089 41,071 CIL Text classification BBC News, Newsgroups, Complaints English 3 tasks
EPI (Wang et al., 2023c) 12,840 3,524 6,917 23,281 CIL Mixed 2 classification tasks AGNews, Amazon Reviews, Yelp Reviews, DBpedia, Yahoo Answers, Web of Science (WOS) English 13 classes
@@
PAGeR (Varshney et al., 2022) 59,754 7,115 15,304 82,173 CIL Intent classification CLINC150, HWU64, BANKING77, Stackoverflow S20, SGD, MWOZ, FewRel English 355 classes
PLE (Li et al., 2022b) 4,669 4,650 31,642 40,961 CIL Intent classification CLINC150, ATIS, HWU64, BANKING77, MTOP, SNIPS, LEYZER, MSLU, TOP English 477 classes
@@
ExtendNER (Monaikul et al., 2021) 130,495 22,717 16,077 169,289 CIL Named Entity Recognition CoNLL-03, OntoNotes English 4 classes